Determine the given, formulate the null and alternative hypothesis in words and in

symbols, and the appropriate test statistic.


A company produced ethyl alcohol and claimed to have a mean alcohol content of 70%. A random sample of 80 of ethyl alcohol was take as sample to verify this claim. It was found out that the mean alcohol content is 65% with a standard deviation of 2%. Test the claim at 5% level of significance and assume that the population is normally distributed.


1
Expert's answer
2022-05-09T18:33:51-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=70," alcohol content is 70%

"H_a:\mu\not=70," alcohol content is not 70%

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=79" degrees of freedom, and the critical value for a two-tailed test is "t_c = 1.99045."

The rejection region for this two-tailed test is "R = \{t: |t| > 1.99045\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{65-70}{2/\sqrt{80}}\approx-22.3607"


5. Since it is observed that "|t| =22.3607>1.99045=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=79" degrees of freedom, "t=-22.3607" is "p=0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 70, at the "\alpha = 0.05" significance level.